# LLMs Capture Urban Science but Oversimplify Complexity

### v2

@@ 4.4 Jane Jacobs’ urban vitality
Fundings
This work was supported by the National Natural Science Foundation of China [Grant No. 62394331] and [Grant No. 62394335].
@@
We have provided information on all publicly available data used in our analysis in the Methods.
@@
The code to replicate the results in the paper will be available via Figshare at ”https://doi.org/10.6084/m9.figshare.28910084”.
@@
Supplementary Note. Models Selection
@@
As evidenced in Figure, GPT-4o and DeepSeek-V3 demonstrated superior alignment with real-world data among the five evaluated models. Under the JSD metric, these two models exhibited aggregate mean divergence values of 0.115 (R-G) and 0.116 (R-GC), representing reductions of $41.9\%$, $23.2\%$, and $1\%$ compared to Claude-3.5, Gemini-2.0, and ChatGLM-4, respectively.
Real data (R) comprises statistical records of population, GDP, and road network length across 685 Chinese municipal administrative regions. Each iteration involved random sampling of 100 cities, with this process repeated 100 times to ensure statistical robustness. Generated data (G) were generated by corresponding models, producing 100 city-level datasets per iteration over 100 cycles. The distinction between R-G origin and R-GC lies in the inclusion of the geographical cue ’China’ in the latter’s generation prompts. Pairwise Jensen-Shannon Divergence (JSD) analyses were conducted to quantify discrepancies among three datasets: real data (R), baseline generated data (G), and generated data with geographical prompt ’China’ (GC).
@@
We evaluated the ability of large language models (LLMs) to generate human height data. Given that height distributions in the real world follow well-documented statistical patterns, such as normal distributions with known mean and standard deviation values for different populations, this task serves as a baseline assessment of whether LLM-generated data aligns with real-world human height distributions. We conducted ten independent trials for each LLM to generate male and female height datasets. The generated data were then compared against the global average height distributions for males (mean = 178.4 cm, SD = 5.6 cm) and females (mean = 164.7 cm, SD = 4.7 cm), as reported by Silverman (2022). LLM-generated height distributions generally conformed to a normal pattern, though accuracy varied notably across models. Claude-3.5 consistently delivered the most realistic estimates, closely matching both the mean and standard deviation of true height data across genders. GPT-4o followed closely, offering stable performance with slightly conservative estimates. In contrast, DeepSeek-V3 exhibited the largest deviations, producing overly dispersed distributions that failed to capture real-world variance. Gender-based differences were also observed. For male height, GPT-4o achieved the best fit, with the highest R² (0.95) and near-identical distribution characteristics, while Claude-3.5 and ChatGLM-4 also performed well. For female height, Claude-3.5 was the most accurate overall, with the smallest deviation from the true mean and standard deviation, followed by ChatGLM-4, which demonstrated the best curve fit (R² = 0.95).
@@
We evaluated LLMs’ ability to generate individual-level wealth data and observed that the resulting distributions generally conformed to a Pareto pattern, capturing the right-skewed nature of wealth concentration. However, the estimated Pareto coefficients were consistently lower than the empirical benchmark of  1.5, as reported by Atkinson [atkinson2014], suggesting that LLMs tend to understate the extent of real-world wealth inequality. Despite this, the log-log linear fits between rank and wealth yielded relatively high R² values, indicating that LLMs can reproduce the structural form of the Pareto distribution. However, high R² does not imply accurate calibration of inequality levels, as the slope of the fitted line (i.e., the Pareto coefficient) remained biased toward more equal distributions. Among the tested models, Gemini-2.0 generated the most realistic distribution, producing both Pareto coefficients and tail behaviors closest to empirical values, thereby better reflecting the heavy-tailed nature of actual wealth distributions. Real-world wealth decile data used for comparison were obtained from the World Inequality Database [wid2025].
Figure S2: Main LLMs tests: Human height distribution.
Figure S3: Main LLMs tests: Individuals’ wealth data distribution.
@@
Figure S5: Main LLMs tests of Distance decay theory: In terms of population density.
Figure S6: Main LLMs tests of Distance decay theory: In terms of urban land density.
Figure S7: Main LLMs tests of Zipf’s Law between cities.
Figure S8: Main LLMs tests of Urban Scaling Law in cities: infrastructure variables.
Figure S9: Main LLMs tests of Urban Scaling Law in cities: socio-economic variables.
@@
Figure S11: Main LLMs tests of Jane Jacobs’ urban vitality theory. E = effective significant positive slope, P ¡ 0.05 and slope ¿ 0.

### v3

@@ 4.4 Post-hoc calibration procedure for data distribution
We have provided information on all publicly available data used in our analysis in the Methods. Please refer to Table S3 for all prompts in this study.
@@
The code to replicate the results in the paper will be available via Figshare.
@@
Supplementary Note 1. Models selection
@@
Supplementary Note 2. Detailed explanations for selected theories
Urban scaling law
Urban distance decay
The term distance decay is used to describe the diminishing intensity of interaction between two locations as the distance between them increases. It refers to the phenomenon where the strength of connections—especially in economic activities—tends to weaken with greater spatial separation [47]. At its core, the distance decay law posits that interactions between geographic entities are inversely related to distance; under otherwise equal conditions, such interactions typically follow an inverse-square relationship with distance. As a concept that reflects the decline of influence over space, distance decay is closely related to Tobler’s First Law of Geography [48]: ’Everything is related to everything else, but near things are more related than distant things’. In urban geography, cities and towns often have clearly defined centers, and most urban attributes—such as population density, land value, and the quality or quantity of services—tend to decrease with increasing distance from the center [47]. Due to spatial agglomeration, urban spatial structures frequently exhibit a pronounced ’center-periphery’ pattern, where the intensity of urban elements decays with increasing distance from the city core. The most typical manifestation of this is the spatial decay of population density. Common quantitative models include the negative exponential model [49] and the inverse power model [50]. For urban land use density, the inverse S-function model has also been proposed to characterize its spatial decay [51].
Jane Jacobs’ urban vitality
Urban vitality theory is a classic urban theory put forward by Jacobs [35]. She pointed out that urban vitality often depends on four key elements: concentration (population density), functional diversity (architectural function mixture), contact opportunity (short block), and need for aged buildings (proportion of old buildings). These factors together constitute the foundation of urban vitality, among which population density and architectural function mixing are the core driving forces, while short blocks and old buildings enhance urban vitality by increasing contact opportunities and reducing operating costs. However, in recent years, there have been disputes about Jacobs’ views on high-rise buildings, and some studies believe that it has a positive impact on urban vitality [52]. Although Jacobs’ theory is widely accepted, empirical tests remain limited and inconsistent. The reasons for uncertainty are the robustness of the research design and the interpretability of the theory to cross-geographical areas.
Urban perception through images
The AI4US framework extends its evaluation beyond statistical regularities to the domain of urban visual perception. Urban science is concerned not only with emergent patterns but also with the human-centric experience of the built environment, which is profoundly shaped by its visual characteristics. Understanding the link between physical appearance and human psychological responses, such as feelings of safety or aesthetic pleasure, is fundamental to urban design and planning.
The study of urban perception has evolved significantly over the past half-century. It originated from foundational, qualitative approaches, such as Kevin Lynch’s work on the urban image [62], which relied on manual surveys and expert analysis to decode how people form mental maps of their cities. The subsequent integration of photography and GIS allowed researchers to begin spatializing these subjective perceptions. A paradigm shift occurred with the rise of computational methods and large-scale crowdsourcing, exemplified by the MIT Place Pulse project [37, 53, 54, 55, 56]. This initiative collected millions of human judgments on street view imagery (SVI), creating a global dataset that links images to perceptual scores. This enabled the next stage: using machine learning to automate perception analysis at an unprecedented scale, and leveraging Explainable AI (XAI) to quantify which visual elements correlate with specific perceptions [57]. The current era of generative multimodal models marks a new frontier, allowing for end-to-end experiments where scenes can be generated, programmatically modified, and re-evaluated to test causal hypotheses about perception.
Supplementary Note 3. Post-hoc calibration procedure for data distribution
To overcome the systematic bias and lack of diversity observed when using GenAI to synthesize data, it is necessary to calibrate the larger dataset generated by GenAI by utilizing a small portion of real empirical data. This challenge of aligning a generated (’source’) distribution with an empirical (’target’) distribution is a fundamental problem in machine learning [58]. It is conceptually analogous to domain adaptation in computer vision, where synthetic data is transformed to match real-world image characteristics [59], or to batch effect correction in bioinformatics, where data from different experiments are normalized to a common distribution [60]. In these contexts, the goal is to learn a mapping that corrects for distributional shifts while preserving the core data structure.
Within the specific context of GenAI-based data synthesis, prior research has largely focused on enhancing output fidelity through sophisticated prompt engineering or agent-based architectures [30, 31]. Post-processing, when applied, has typically involved direct, rule-based adjustments. These include enforcing strict output formats and value ranges to ensure data validity at the point of generation [30], or subsequently grounding the GenAI’s conceptual outputs (e.g., a ”trip to the office”) in physical space by mapping them to specific geographic coordinates based on real-world constraints like a city’s road network [26]. While these methods improve the plausibility of individual data points, they may not correct for systematic biases in the overall data distribution.
To address this need, we selected an approach grounded in Optimal Transport (OT) [58, 61] theory for its distinct advantages over simpler alternatives. Direct probability density estimation methods (e.g., Kernel Density Estimation) often struggle with the curse of dimensionality in multi-attribute urban data. Simpler filtering methods, such as classifier-based selection, can discard a significant portion of the generated samples and may fail to preserve the latent structure of the data. In contrast, OT offers a more powerful solution by transforming the entire generated dataset rather than merely selecting from it, providing a robust method for distributional alignment that is well-suited to this task.
Table S1: A Multi-Scale Framework for Evaluating GenAI models in Urban Science.
Scale Theory Significance in Urban Science
Macro (City system) Urban scaling law Explains universal, size-based regularities of cities as complex systems, central to understanding urban growth and efficiency.
Meso (Internal city structure) Urban distance decay Represents the foundational role of spatial friction and accessibility in shaping internal urban structure and organization.
Micro (Block) Urban vitality Provides a human-centric perspective linking the micro-design of the built environment to social and economic outcomes; a cornerstone of urban design.
Table S2: Benchmark for Urban Perception Indicators based on Visual Elements. Summary of regression coefficients ($\beta$) for all unique top-10 visual objects influencing six urban perceptions. The significance of each coefficient is indicated by stars. Adapted from the analysis in Zhang et al. (2018) [57]. The top 10 objects that positively/negatively contributed to each of the 6 perception types are shown. Beta Coefficient (*p¡0.1,**p¡0.05,***p¡0.01).
Visual Object Safe Lively Beautiful Wealthy Depressing Boring
Bridge - -0.0144*** -0.0120** - 0.0178*** 0.0076
Building -0.1163*** -0.0625*** -0.0689*** - 0.0483*** -
Car 0.0600*** 0.0887*** - 0.0295*** -0.0243*** -0.0329***
Earth - - - -0.0174*** - -
Fence -0.0327*** - -0.0201*** - 0.0198*** 0.0144***
Field - -0.0543*** - -0.0248*** - 0.0130**
Floor - - -0.0172*** -0.0219*** - -
Grass 0.0711*** - 0.0465*** 0.0272*** -0.0204*** -
Hill - - - - - 0.0105*
House - - - - -0.0123** -
Path 0.0154*** - 0.0126*** 0.0115** - -
Pole - - - - - -0.008
Road 0.0395*** 0.0228*** - 0.0122* - -
Signboard - - - - - 0.0173***
Sidewalk 0.0821*** 0.0596*** - - -0.0350*** -0.0320***
Sky -0.2068*** -0.2389*** -0.1179*** -0.1142*** 0.1063*** 0.0969***
Tree -0.0392*** -0.1055*** 0.0980*** 0.0244*** -0.0610*** -
Truck - - -0.0125*** - - -
Wall -0.0764*** -0.0828*** -0.0458*** -0.0264*** 0.0389*** 0.0382***
Water - - 0.0070* - - -
Table S3: Prompts and specific descriptions in this study.
Prompt Description
Urban scaling law Generate the dataset containing 100 cities in a country. Format the output as: CityName1, Population1, Infrastructure volume1, GDP1 CityName2, Population2, Infrastructure volume2, GDP2 Output exactly 100 lines without any additional text or explanations. - Infrastructure volume: Total road miles - GDP: all Gross Domestic Product of the city in one year
Urban vitality theory – generation Generate attributes for 100 blocks in a city. Format the output as: Block Name1, Population Density1, Building Mix Index1, Short Block1, Aged Building1, Tall Building1 Output exactly 100 lines without any additional text or explanations. - Population Density: People per square meter. - Building Mix Index: Entropy index based on building use. Value 0 to 1. - Short Block: Total number of street intersections divided by the area of the block. - Aged Building: The rate of old buildings on each block. Value 0 to 1. - Tall Building: The average building heights on each block.
Urban vitality theory – evaluation You are a resident of a city. Assign a livability score between 0 and 1 for each block, 0 represents the least livable and 1 represents the most livable. Evaluate each block based on the following attributes: - Block name. - Population Density: People per square meter. - Building Mix Index: Entropy index based on building use. - Short Block: Total number of street intersections of the block. - Aged Building: The rate of old buildings on each block. - Tall Building: The average building heights on each block.
Streetscape Generation (Phase 1 Base Template) ”A photorealistic, eye-level street view of {scene_description} from a random place on earth. The screen should be completely filled with the generated image, without white borders. Try to approach the real scene captured by the car (A specialized vehicle equipped with professional camera equipment for capturing street view images) as closely as possible, without any retouching or beautification. The camera perspective is a forward-facing view, looking straight down the road. It should be a head-up view without any perspective distortion.”
Causal Intervention (Phase 3 Base Template) Realistically edit the image to {intervention_description}. Keep everything else in the image the same.
@@
As evidenced in Figure S1, GPT-4o and DeepSeek-V3 demonstrated superior alignment with real-world data among the five evaluated models. Under the JSD metric, these two models exhibited aggregate mean divergence values of 0.115 (R-G) and 0.116 (R-GC), representing reductions of 41.9%, 23.2%, and 1% compared to Claude-3.5, Gemini-2.0, and ChatGLM-4, respectively. Real data (R) comprises statistical records of population, GDP, and road network length across 685 Chinese municipal administrative regions. Each iteration involved random sampling of 100 cities, with this process repeated 100 times to ensure statistical robustness. Generated data (G) were generated by corresponding models, producing 100 city-level datasets per iteration over 100 cycles. The distinction between R-G origin and R-GC lies in the inclusion of the geographical cue ’China’ in the latter’s generation prompts. Pairwise Jensen-Shannon Divergence (JSD) analyses were conducted to quantify discrepancies among three datasets: real data (R), baseline generated data (G), and generated data with geographical prompt ’China’ (GC).
@@
In this supplementary experiment, we evaluated the ability of GenAI models to generate human height data. Given that height distributions in the real world follow well-documented statistical patterns, such as normal distributions with known mean and standard deviation values for different populations, this task serves as a baseline assessment of whether generated data aligns with real-world human height distributions. We conducted ten independent trials for each GenAI model to generate male and female height datasets. The generated data were then compared against the global average height distributions for males (mean = 178.4 cm, SD = 5.6 cm) and females (mean = 164.7 cm, SD = 4.7 cm), as reported by Silverman (2022). Generated height distributions generally conformed to a normal pattern, though accuracy varied notably across models. Claude-3.5 consistently delivered the most realistic estimates, closely matching both the mean and standard deviation of true height data across genders. GPT-4o followed closely, offering stable performance with slightly conservative estimates. In contrast, DeepSeek-V3 exhibited the largest deviations, producing overly dispersed distributions that failed to capture real-world variance. Gender-based differences were also observed. For male height, GPT-4o achieved the best fit, with the highest R² (0.95) and near-identical distribution characteristics, while Claude-3.5 and ChatGLM-4 also performed well. For female height, Claude-3.5 was the most accurate overall, with the smallest deviation from the true mean and standard deviation, followed by ChatGLM-4, which demonstrated the best curve fit (R² = 0.95).
Figure S2: Main GenAI models tests: Human height distribution.
@@
We evaluated GenAI’s ability to generate individual-level wealth data and observed that the resulting distributions generally conformed to a Pareto pattern, capturing the right-skewed nature of wealth concentration. However, the estimated Pareto coefficients were consistently lower than the empirical benchmark of  1.5, as reported by Atkinson and Piketty, suggesting that GenAI models tend to understate the extent of real-world wealth inequality. Despite this, the log-log linear fits between rank and wealth yielded relatively high R² values, indicating that GenAI models can reproduce the structural form of the Pareto distribution. However, high R² does not imply accurate calibration of inequality levels, as the slope of the fitted line (i.e., the Pareto coefficient) remained biased toward more equal distributions. Among the tested models, Gemini-2.0 generated the most realistic distribution, producing both Pareto coefficients and tail behaviors closest to empirical values, thereby better reflecting the heavy-tailed nature of actual wealth distributions. Real-world wealth decile data used for comparison were obtained from the World Inequality Database.
Figure S3: Main GenAI models tests: Individuals’ wealth data distribution.
@@
Figure S5: Main GenAI models tests of Distance decay theory: In terms of population density.
Figure S6: Main GenAI models tests of Distance decay theory: In terms of urban land density.
Figure S7: Main GenAI models tests of Zipf’s Law between cities.
Figure S8: Main GenAI models tests of Urban Scaling Law in cities: infrastructure variables.
Figure S9: Main GenAI models tests of Urban Scaling Law in cities: socio-economic variables.
@@
Figure S11: Main GenAI models tests of Jane Jacobs’ urban vitality theory. E = effective significant positive slope, P ¡ 0.05 and slope ¿ 0.
Figure S12: Enhancing the effectiveness of GenAI as a data synthesizer based on optimal transport. The top panels display the mean relative frequency distributions, while the bottom panels provide a quantitative evaluation using three key performance metrics. All results are aggregated from 100 independent experimental runs. In each run, a large synthetic dataset (1,000 data points) is generated by each method, calibrated against a new random sample of 100 real cities.
Figure S13: Urban streetscape image realism evaluation.